Rudd Clothiers is a small company that manufactures tall-men’s suits. The company has used a standard cost accounting system. In May 2017, 10,600 suits were produced. The following standard and actual cost data applied to the month of May when normal capacity was 16,000 direct labor hours. All materials purchased were used.

Cost Element

Standard (per unit)

Actual

Direct materials 10 yards at $4.90 per yard $501,020 for 106,600 yards ($4.70 per yard)
Direct labor 1.20 hours at $13.00 per hour $182,508 for 13,620 hours ($13.40 per hour)
Overhead 1.20 hours at $6.40 per hour (fixed $3.60; variable $2.80) $49,200 fixed overhead $36,500 variable overhead


Overhead is applied on the basis of direct labor hours. At normal capacity, budgeted fixed overhead costs were $57,600, and budgeted variable overhead was $44,800.

(a)

Compute the total, price, and quantity variances for (1) materials and (2) labor. (Round answers to 0 decimal places, e.g. 125.)

(1) Total materials variance
Materials price variance
Materials quantity variance

Entry field with incorrect answer

(2) Total labor variance

Entry field with incorrect answer

Labor price variance
Labor quantity variance


(b)

Compute the total overhead variance.

Total overhead variance

Entry field with incorrect answer
Answer #1

a. 1. Total material variance = 10600*10*4.90 - 501020 = 18380 favorable

Material price variance = (4.90-4.70)*106600 = 21320 favorable

Material quantity variance = (10600*10 - 106600)*4.90 = 2940 unfavorable

2. Total labor variance = 10600*1.2*13 - 182508 = 17148 unfavorable

Labor price variance = (13-13.40)*13620 = 5448 unfavorable

Labor quantity variance = (10600*1.2-13620)*13 = 11700 unfavorable

b. Total overhead variance = 16000*3.60 + 10600*1.20*2.80 - 49200+36500 = 7516 favorable